OpenStudy (anonymous):

The stopping distance s of a car varies directly as the square of its speed v. If a car traveling 30 mph requires 45 ft to stop, find the stopping distance for a car traveling 65 mph.

OpenStudy (anonymous):

\[s=20(65)^2\]

OpenStudy (radar):

s=kv^2 45=k30^2 45=k900 k=45/900=1/20 s=(1/20)65^2 s=4225/20=211.25 ft.

OpenStudy (anonymous):

Thank You!!

OpenStudy (radar):

@ebbflo, is that your dog in the logo?

OpenStudy (radar):

I believe ebbflo meant (1/20)(65)^2

OpenStudy (anonymous):

no that's me:) yeah it is one of my dogs

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

sorry

OpenStudy (radar):

Beautiful animal. As you may see my logo has my dog, a boston terrier.